OpenStudy (anonymous):

Solve 7 + x - 15 = -2. 5 6 10 20

OpenStudy (anonymous):

\[7 + x - 15 = -2\]\[7 + x - 15 +15= -2+15\]\[7 + x = 13\]\[7 + x -7= 13-7\]\[x=~?~~you~~tell~~me.\]

OpenStudy (anonymous):

idk thats why i asked

OpenStudy (anonymous):

I did all the steps. You just have to do the last hit.

OpenStudy (anonymous):

\(x=6\)

OpenStudy (anonymous):

Yeah, how great we are at giving the answers.... not even thinking about anything.....

OpenStudy (anonymous):

Solve for y. 7y - 6y - 10 = 13 y = -23 y = -3 y = 3 y = 23

OpenStudy (anonymous):

7y-6y-10=13 y-10=13 y=23

OpenStudy (anonymous):

Which of the following equations has the solution set {0}? -14m - 7m + 1 = -7m + 1 -14m + 7m + 1 = -7m + 2 14m - 7m - 1 = -7m + 1

OpenStudy (anonymous):

\( 7y-6y-10=13\) \(\Leftrightarrow y=13+10\) \(\Leftrightarrow y=23\)

OpenStudy (anonymous):

The equation has the solution set {0} is \(-14m - 7m + 1 = -7m +1\) Because: If \(m=0\), them \( -14\times 0 - 7\times 0 + 1 = -7\times 0 + 1\) \(\Rightarrow 1=1\) (always satisfy)